OpenStudy (anonymous):

Write each function in vertex form, and identify its vertex. Plz show steps! f(x)=x^2+24x+138

OpenStudy (anonymous):

for the vertex we need a pair (x,y) to find x we have a formula x=-b+2a

OpenStudy (anonymous):

lets identify a and b a is the coefficient of x^2 thats is 1 and be is the coefficient of x that is 24

OpenStudy (anonymous):

x=-24/2(1) x=-24/2 x=-12

OpenStudy (anonymous):

then to find y we plug -12 in the equation (-12)^2+24(-12)+138

OpenStudy (anonymous):

can u guess the answer?

OpenStudy (anonymous):

Yeah. Y would be -6

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so your vertex will be (-12,-6)

OpenStudy (anonymous):

Oh, great thanks!